# Variational quantum state preparation for quantum-enhanced metrology in noisy systems

### v1

@@ Abstract
We investigate optimized quantum state preparation for quantum metrology applications in noisy environments. We simulate a low-depth variational quantum circuit (VQC) composed of a sequence of global rotations and entangling operations applied to a chain of qubits that are subject to dephasing noise. The parameters controlling the VQC are numerically optimized to maximize the quantum Fisher information, which characterizes the ultimate metrological sensitivity of a quantum state, with respect to a global rotation. We find that regardless of the details of the entangling operation implemented in the VQC, the optimal quantum states can be broadly classified into a trio of qualitative regimes – cat-like, squeezed-like, and product states – associated with different dephasing rates. Our findings are relevant for designing optimal state-preparation strategies for next-generation quantum sensors exploiting entanglement, such as time and frequency standards and magnetometers, aimed at achieving state-of-the-art performance in the presence of noise and decoherence.
@@ I Introduction
Quantum sensing holds great promise amongst emerging quantum technologies that exploit entanglement and quantum correlations as the closest to achieving a near-term, substantive quantum advantage over state-of-the-art classical devices. Prominent successes include LIGO, which has increased its detection volume with frequency-dependent squeezed light LIGO O4 Detector Collaboration et al. (2023), and the use of squeezed quantum states in searches for dark-matter candidates Jewell et al. (2023). Similarly, optical frequency metrology experiments are at the precipice of benefiting from quantum entanglement and correlations to push measurement precision beyond the limits of quantum projection noise Colombo et al. (2022); Robinson et al. (2024); Bothwell et al. (2022); Pedrozo-Peñafiel et al. (2020).
@@
Progress towards the paradigm of quantum-enhanced sensing is being fueled through rapid improvements in technical capabilities, such as the realization of fully controllable intermediate scale quantum hardware featuring, e.g., 10s to even 1000s of high-quality qubits Ebadi et al. (2021); Manetsch et al. (2024). These systems exhibit long coherence times, an exquisite degree of coherent single-qubit control, and innovative methods for many-body control (i.e., entangling operations) via natural or engineered interactions Britton et al. (2012); Katz et al. (2023); Bornet et al. (2023a); Miller et al. (2024); Zhang et al. (2023).
@@
A key challenge for current research is to establish robust metrological schemes that exhaust the potential of these hardware platforms in the presence of technical imperfections and decoherence. In the absence of these, optimal entangled quantum states for metrology are often exactly known, and can be obtained by analytically maximizing the quantum Fisher information $F_{Q}$, which bounds from below the achievable variance $(\Delta\theta)^{2}$ in an estimation parameter $\theta$ via the quantum Cramer-Rao bound, $(\Delta\theta)^{2}\geq 1/F_{Q}$ Braunstein and Caves (1994); Skotiniotis et al. (2015); Liu et al. (2019). In the context of sensing spin rotations, there exist a diversity of protocols, including non-equilibrium and quasi-adiabatic dynamics Lee (2006); Leibfried et al. (2005); Micheli et al. (2003); Nolan and Haine (2018), to prepare these ideal entangled states as well as optimize signal acquisition and readout to infer $\theta$ Leibfried et al. (2004); Haine (2018), and which can be tailored for specific experimental platforms. However, these optimal quantum states – corresponding to cat-like macroscopic superposition states – are catastrophically susceptible to decoherence Huelga et al. (1997); Pezzè et al. (2018), which is ever-present in real-world quantum hardware. It remains an open challenge to identify and prepare states that maximize quantum Fisher information in the presence of decoherence.
@@
In this work, we perform a preliminary study of optimal state preparation for metrology in a noisy quantum environment. Our work is motivated by an effort to bring to bear existing resources and coherent control to maximize the potential of near-term quantum devices to demonstrate a state-of-the-art quantum advantage in sensing tasks, as opposed to proof-of-principle realizations. To make our task concrete, we numerically optimize a simple, low-depth variational quantum circuit (VQCs, see Fig. 1) for state preparation in an effort to maximize the resulting quantum Fisher information. Motivated by current physical platforms for quantum computation and simulation, we include a range of entangling gates in our VQC based on a variety of natural or engineered two-qubit interactions. We seek to maximize the QFI of the state generated by the VQC for a range of experimentally relevant noise models. We find that the optimal states can typically be grouped into a number of generic regimes – such as cat-like and squeezed-like quantum states – regardless of the nature of the entangling interactions, with some caveats associated with interaction range and system size. While our calculations are limited to relatively small systems (because the QFI for mixed quantum states is computationally expensive to obtain), we comment on implications for larger ensembles of qubits.
@@
Our paper is structured as follows. In Sec. II we outline our basic VQC, including fundamental entangling operations and sources of decoherence. We present our results for the optimal QFI obtained with these VQCs in Sec. III, which includes a detailed analysis of the structure of the obtained optimal quantum states as well as a discussion of their dependence on system size and interaction range. We summarize our findings and present an outlook in Sec. IV.
@@ II Variational quantum state preparation
As a basis for our investigation, we consider a low-depth VQC to prepare a quantum state of $N$ qubits to be used for sensing global qubit rotations. The applications for such states are diverse and include, e.g., frequency metrology and magnetometry. The VQC consists of two key pieces: (i) a set of global rotations occurring during the initial and final stages to prepare and reorient the qubits for sensing, and (ii) an intermediate stage where the qubits interact to create entanglement. In more detail (see Fig. 1), the VQC sequence begins by preparing an ensemble of $N$ qubits in a pure product state |$\ket{\psi_{0}}=\ket{\downarrow}^{\otimes N}$. The initial and final rotations are parameterized by $\hat{R}_{\theta_{j}}^{\varphi_{j}}=\prod_{k=1}^{N}e^{-i\theta_{j}\hat{\sigma}^{\varphi_{j}}_{k}/2}$ where $\hat{\sigma}_{k}^{\varphi_{j}}=\hat{\sigma}^{x}_{k}\cos{(\varphi_{j})}+\hat{\sigma}_{k}^{y}\sin{(\varphi_{j})}$ are Pauli matrices for the $k$th qubit in a rotated basis where $\theta_{j}$ and $\varphi_{j}$ define the rotation angle and axis for the initial ($j=1$) or final ($j=2$) rotation. The intermediate entangling step is generated by the gate:
@@
where the Hamiltonian $\hat{H}_{\mathrm{I}}$ describes the interaction of the qubits with characteristic coupling strength $\chi$ (see below). We parameterize the “strength” of the entangling gate through the quantity $\theta_{\mathrm{I}}=\chi t_{\mathrm{I}}$ where $t_{\mathrm{I}}$ is the duration of the applied gate. Note that this preliminary description of the entangling gate as a unitary operation ignores the role of decoherence, but we will address this in the following subsections.
In total, the VQC is characterized by a set of five parameters $\mathbf{x}=(\theta_{1},\varphi_{1},\theta_{I},\theta_{2},\varphi_{2})$ and additionally through the choice of entangling Hamiltonian, $\hat{H}_{\mathrm{I}}$. We denote the output state of the VQC in terms of these parameters via the density matrix $\hat{\rho}(\mathbf{x})$.
@@ II.1 Entangling gates
We study a variety of Hamiltonians $\hat{H}_{\mathrm{I}}$ for the entangling gate, motivated by both conceptual and practical considerations. Firstly, to benchmark our findings we explore one-axis (OAT) and two-axis (TAT) twisting Hamiltonians featuring infinite-range pairwise interactions between qubits Kitagawa and Ueda (1993). The generation of entangled states for metrology using the unitary dynamics generated by these Hamiltonians, such as spin-squeezed and Greenberger-Horne-Zeilinger (GHZ) states, is well established Kitagawa and Ueda (1993); Tóth and Apellaniz (2014); Pezzè et al. (2018). In particular, OAT dynamics have been realized and exploited in a diversity of experimental platforms including trapped ions Bohnet et al. (2016), Bose-Einstein condensates Muessel et al. (2014) and cavity-QED Leroux et al. (2010); Hosten et al. (2016); Norcia et al. (2018); Braverman et al. (2019), and there has been early work proposing He et al. (2019) and even demonstrating Miller et al. (2024); Luo et al. (2024) TAT interactions. In addition, motivated by current efforts to generate entangled states in quantum simulation platforms featuring finite-range pairwise couplings between qubits Perlin et al. (2020); Young et al. (2023); Eckner et al. (2023); Franke et al. (2023); Bornet et al. (2023b), we also study Ising, XX, and f-TAT (finite-range TAT) spin chain models with power law interactions. The Hamiltonians for these models are
@@
$\hat{H}_{\mathrm{XX}}$ $=\sum_{i\neq j}{\chi_{ij}(\hat{\sigma}_{i}^{x}\hat{\sigma}_{j}^{x}+\hat{\sigma}_{i}^{y}\hat{\sigma}_{j}^{y}}),$ (3)
@@
Here $i$ and $j$ index individual qubits at positions $r_{i}$ and $r_{j}$, and the couplings $\chi_{ij}=\chi/\mathcal{N}_{\alpha}|r_{i}-r_{j}|^{\alpha}$, where $\chi$ sets the strength of interactions and $\mathcal{N}_{\alpha}=(N-1)^{-1}\sum_{i\neq j}|r_{i}-r_{j}|^{-\alpha}$ is a normalization factor to ensure that the Hamiltonian is extensive Uhlenbeck et al. (1963); Defenu (2021) and simplify comparison of the models as a function of decoherence strength (see below). Throughout this manuscript we set $\hbar=1$ and, for simplicity, we focus on qubits in a one-dimensional geometry that can be realized using, e.g., arrays of neutral atoms or trapped ions.
We will restrict our focus to interactions featuring power-law exponents $\alpha=0,...,6$, motivated by the range featured in prominent quantum science platforms such as Rydberg atoms ($\alpha=3,6$), trapped ions ($\alpha\in[0,3]$), molecules ($\alpha=3$) and light-matter systems ($\alpha=0$). The limiting case of $\alpha=0$ corresponds to all-to-all interactions, for which the Ising and XX models reduce (up to a constant of motion) to the OAT Hamiltonian $\propto\hat{S}_{z}^{2}$, while f-TAT resolves to the TAT Hamiltonian $\propto(\hat{S}_{x}\hat{S}_{y}+\hat{S}_{y}\hat{S}_{x})$.
II.2 Dephasing noise
Motivated by the implementation of the above VQC in real-world hardware, we incorporate into our description decoherence due to, e.g., stray background fields and spurious couplings to the environment. However, in our treatment we will make the simplifying assumption that decoherence only occurs during the application of the entangling gate, with the dynamics of the initial and final rotations taken to be sufficiently rapid that they may be approximated as unitary processes.
The inclusion of decoherence and dissipation means that the unitary representation of the entangling gate given by Eq. (1) is inappropriate. Instead, we model this stage of the VQC using a Linblad master equation,
$\frac{d\hat{\rho}}{dt}=-i[\hat{H}_{\mathrm{I}},\hat{\rho}]+\sum_{\nu,j}\gamma_{\nu}\left(\hat{L}^{\nu}_{j}\hat{\rho}\hat{L}^{\nu\dagger}_{j}-\frac{1}{2}\{\hat{L}^{\nu\dagger}_{j}\hat{L}^{\nu}_{j},\hat{\rho}\}\right).$ (5)
where $\hat{H}_{\mathrm{I}}$ is the chosen entangling Hamiltonian [Eqs. (2)-(4)], $\nu\in\set{x,\;y,\;z}$ indexes an axis for dephasing at rate $\gamma_{\nu}$, and $j\in\set{1,2,\cdots,N}$ indexes a qubit addressed by the jump operator $\hat{L}^{\nu}_{j}=\hat{\sigma}^{\nu}_{k}/2$. In this work, we will assume that the dephasing is isotropic, i.e., acts with equal rates $\gamma_{x}=\gamma_{y}=\gamma_{z}=\gamma$. This assumption avoids bias in our results due to, e.g., robustness of the entangling dynamics for specific representations of the Hamiltonians in Eqs. (2)-(4Britton et al. (2012); Jurcevic et al. (2014). In future work Zuñiga Castro et al. (ming), we will consider the generic case of anisotropic dephasing (unequal $\gamma_{x},\gamma_{y},\gamma_{z}$) as well as decoherence due to spontaneous emission and absorption.
@@ II.3 Optimization of the quantum Fisher information
We seek to identify optimal state preparation protocols, within the constraints of the VQC defined above, for metrology of global spin rotations. The quantity capturing the suitability of a generic quantum state $\hat{\rho}$ for this task is the quantum Fisher information Pang and Brun (2014)
@@
Here |$\ket{k}$ and $\lambda_{k}$ are, respectively, the eigenvectors and corresponding eigenvalues of the density matrix prepared by the VQC, i.e., $\hat{\rho}=\sum_{k}\lambda_{k}\ket{k}\!\bra{k}$ |, and $\hat{G}$ is the generator of a unitary transformation $\hat{U}_{\phi}=e^{-i\phi\hat{G}}$ encoding a classical parameter $\phi$. The QFI characterizes the uncertainty with which the classical parameter can be estimated through the quantum Cramer-Rao bound (QCRB), $(\Delta\phi)^{2}\geq 1/F_{Q}[\hat{\rho}(\mathbf{x});\hat{G}]$. Without loss of generality, throughout this work we set $\hat{G}=\hat{S}_{z}=\sum_{j}\hat{\sigma}^{z}_{j}/2$.
@@
In the following section, we instead investigate optimal preparation strategies, using the VQC discussed previously, that take into account the effects of decoherence. For a fixed decoherence strength $\gamma$, we obtain the state $\hat{\rho}(\mathbf{x})$ generated by a given VQC with parameters $\mathbf{x}$. We numerically integrate the master equation (5) for a given choice of entangling Hamiltonian $\hat{H}_{\mathrm{I}}$ and associated global rotations. The optimal preparation strategies defined by parameters $\mathbf{x}_{\mathrm{opt}}$ are obtained by numerically optimizing over $\mathbf{x}$ to maximize the QFI of Eq. (6). Throughout this work we use derivative-free algorithms such as the Nelder-Mead method in SciPy Virtanen et al. (2020) to obtain $\mathbf{x}_{\mathrm{opt}}$. The code package used to simulate the VQC is available at Ref. Colussi et al. (2024).
III Results

### v2

@@ Abstract
We investigate optimized quantum state preparation for quantum metrology applications in noisy environments. Using the QFI-Opt package, we simulate a low-depth variational quantum circuit (VQC) composed of a sequence of global rotations and entangling operations applied to a chain of qubits that are subject to dephasing noise. The parameters controlling the VQC are numerically optimized to maximize the quantum Fisher information, which characterizes the ultimate metrological sensitivity of a quantum state with respect to a global rotation. We find that regardless of the details of the entangling operation implemented in the VQC, the optimal quantum states can be broadly classified into a trio of qualitative regimes—cat-like, squeezed-like, and product states— associated with different dephasing rates. Our findings are relevant for designing optimal state-preparation strategies for next-generation quantum sensors exploiting entanglement, such as time and frequency standards and magnetometers, aimed at achieving state-of-the-art performance in the presence of noise and decoherence.
@@ I Introduction
Quantum sensing holds great promise among emerging technologies that exploit quantum phenomena such as entanglement to achieve a near-term, substantive quantum advantage over state-of-the-art classical devices. Prominent successes include LIGO, which has increased its detection volume with frequency-dependent squeezed light LIGO O4 Detector Collaboration et al. (2023), and the use of squeezed quantum states in searches for dark-matter candidates Jewell et al. (2023). Similarly, optical frequency metrology experiments are at the precipice of benefiting from quantum entanglement and correlations to push measurement precision beyond the limits of quantum projection noise Colombo et al. (2022); Robinson et al. (2024); Bothwell et al. (2022); Pedrozo-Peñafiel et al. (2020).
@@
Progress toward the paradigm of quantum-enhanced sensing is being fueled through rapid improvements in technical capabilities, such as the realization of fully controllable intermediate-scale quantum hardware featuring tens to even thousands of high-quality qubits Ebadi et al. (2021); Manetsch et al. (2024). These systems exhibit long coherence times, an exquisite degree of coherent single-qubit control, and innovative methods for many-body control (i.e., entangling operations) via natural or engineered interactions Britton et al. (2012); Katz et al. (2023); Bornet et al. (2023a); Miller et al. (2024); Zhang et al. (2023).
@@
A key challenge for current research is to establish robust metrological schemes that exhaust the potential of these hardware platforms in the presence of technical imperfections and decoherence Jiao et al. (2023); MacLellan et al. (2024); Campbell et al. (2023). In the absence of these, optimal entangled quantum states for metrology can be obtained by analytically maximizing the quantum Fisher information $F_{Q}$, which bounds from below the achievable variance $(\Delta\theta)^{2}$ in an estimation parameter $\theta$ via the quantum Cramer–Rao bound, $(\Delta\theta)^{2}\geq 1/F_{Q}$ Braunstein and Caves (1994); Skotiniotis et al. (2015); Liu et al. (2019). In the context of sensing spin rotations, diverse protocols exist, including non-equilibrium and quasi-adiabatic dynamics Lee (2006); Leibfried et al. (2005); Micheli et al. (2003); Nolan and Haine (2018), to prepare these ideal entangled states as well as optimize signal acquisition and readout to infer $\theta$ Leibfried et al. (2004); Haine (2018); Cao et al. (2024); Finkelstein et al. (2024); Direkci et al. (2024), and which can be tailored for specific experimental platforms. However, these optimal quantum states—corresponding to cat-like macroscopic superposition states—are catastrophically susceptible to decoherence Huelga et al. (1997); Pezzè et al. (2018), which is ever-present in real-world quantum hardware. It remains an open challenge to identify and prepare states that maximize quantum Fisher information (QFI) in the presence of decoherence Fiderer et al. (2019).
@@
In this work we perform a preliminary study of optimal state preparation for metrology in a noisy quantum environment. Our work is motivated by an effort to bring to bear existing resources and coherent control to maximize the potential of near-term quantum devices to demonstrate a state-of-the-art quantum advantage in sensing tasks, as opposed to proof-of-principle realizations. To make our task concrete, we numerically optimize a simple, low-depth variational quantum circuit (VQC; see Fig. 1) for state preparation in an effort to maximize the resulting quantum Fisher information. Motivated by current physical platforms for quantum computation and simulation, we include a range of entangling gates in our VQC based on a variety of natural or engineered two-qubit interactions. We seek to maximize the QFI of the state generated by the VQC for a range of experimentally relevant noise models. We find that the optimal states can typically be grouped into several regimes—such as cat-like and squeezed-like quantum states—regardless of the nature of the entangling interactions, with some caveats associated with interaction range and system size. While our calculations are limited to relatively small systems (because the QFI for mixed quantum states is computationally expensive to obtain), we comment on implications for larger ensembles of qubits.
@@
Our paper is structured as follows. In Sec. II we outline our basic VQC, including fundamental entangling operations and sources of decoherence. We present our results for the optimal QFI obtained with these VQCs in Secs. III and IV, which includes a detailed analysis of the structure of the obtained optimal quantum states as well as a discussion of the dependence on system size and interaction range. In Sec. V we summarize our findings and present suggestions for future research.
@@ II Variational quantum state preparation
As a basis for our investigation we consider a low-depth VQC to prepare a quantum state of $N$ qubits to be used for sensing global qubit rotations. The applications for such states are diverse and include frequency metrology and magnetometry. The VQC consists of two key pieces: (i) a set of global rotations occurring during the initial and final stages to prepare and reorient the qubits for sensing and (ii) an intermediate stage where the qubits interact to create entanglement. In more detail (see Fig. 1), the VQC sequence begins by preparing an ensemble of $N$ qubits in a pure product state |$\ket{\psi_{0}}=\ket{\downarrow}^{\otimes N}$. The initial and final rotations are parameterized by $\hat{R}_{\theta_{j}}^{\varphi_{j}}=\prod_{k=1}^{N}e^{-i\theta_{j}\hat{\sigma}^{\varphi_{j}}_{k}/2}$, where $\hat{\sigma}_{k}^{\varphi_{j}}=\hat{\sigma}^{x}_{k}\cos{(\varphi_{j})}+\hat{\sigma}_{k}^{y}\sin{(\varphi_{j})}$ are Pauli matrices for the $k$th qubit in a rotated basis, where $\theta_{j}$ and $\varphi_{j}$ respectively define the rotation angle and axis for the initial ($j=1$) or final ($j=2$) rotation. The intermediate entangling step is generated by the gate:
@@
where the Hamiltonian $\hat{H}_{\mathrm{I}}$ describes the interaction of the qubits with characteristic coupling strength $\chi$ (see below). We parameterize the “strength” of the entangling gate through the quantity $\theta_{\mathrm{I}}=\chi t_{\mathrm{I}}$, where $t_{\mathrm{I}}$ is the duration of the applied gate. Note that this preliminary description of the entangling gate as a unitary operation ignores the role of decoherence, but we will address this in the following subsections.
In total, the VQC is characterized by a set of five parameters, $\mathbf{x}=(\theta_{1},\varphi_{1},\theta_{I},\theta_{2},\varphi_{2})$, and additionally through the choice of entangling Hamiltonian, $\hat{H}_{\mathrm{I}}$. We denote the output state of the VQC in terms of these parameters via the density matrix $\hat{\rho}(\mathbf{x})$.
@@ II.1 Entangling gates
We study a variety of Hamiltonians $\hat{H}_{\mathrm{I}}$ for the entangling gate, motivated by both conceptual and practical considerations. First, to benchmark our findings, we explore one-axis (OAT) and two-axis (TAT) twisting Hamiltonians featuring infinite-range pairwise interactions between qubits Kitagawa and Ueda (1993). The generation of entangled states for metrology using the unitary dynamics generated by these Hamiltonians, such as spin-squeezed and Greenberger–Horne–Zeilinger (GHZ) states, is well established Kitagawa and Ueda (1993); Tóth and Apellaniz (2014); Pezzè et al. (2018). In particular, OAT dynamics have been realized and exploited in a diversity of experimental platforms including trapped ions Bohnet et al. (2016), Bose–Einstein condensates Muessel et al. (2014), and cavity-QED Leroux et al. (2010); Hosten et al. (2016); Norcia et al. (2018); Braverman et al. (2019); and early work has proposed He et al. (2019) and even demonstrated Miller et al. (2024); Luo et al. (2024) TAT interactions. In addition, motivated by current efforts to generate entangled states in quantum simulation platforms featuring finite-range pairwise couplings between qubits Perlin et al. (2020); Young et al. (2023); Eckner et al. (2023); Franke et al. (2023); Bornet et al. (2023b), we also study Ising, XX, and f-TAT (finite-range TAT) spin chain models with power-law interactions. The Hamiltonians for these models are
@@
$\hat{H}_{\mathrm{XX}}$ $=\sum_{i\neq j}{\chi_{ij}(\hat{\sigma}_{i}^{x}\hat{\sigma}_{j}^{x}+\hat{\sigma}_{i}^{y}\hat{\sigma}_{j}^{y}}),\text{ and}$ (3)
@@
Here $i$ and $j$ index individual qubits at positions $r_{i}$ and $r_{j}$, and the couplings $\chi_{ij}=\chi/\mathcal{N}_{\alpha}|r_{i}-r_{j}|^{\alpha}$, where $\chi$ sets the strength of interactions and $\mathcal{N}_{\alpha}=(N-1)^{-1}\sum_{i\neq j}|r_{i}-r_{j}|^{-\alpha}$ is a normalization factor to ensure that the Hamiltonian is extensive Uhlenbeck et al. (1963); Defenu (2021) and simplify comparison of the models as a function of decoherence strength (see below). Throughout this manuscript we set $\hbar=1$; and, for simplicity, we focus on qubits in a one-dimensional geometry that can be realized using, for example, arrays of neutral atoms or trapped ions.
We will restrict our focus to interactions featuring power-law exponents $\alpha=0,...,6$, motivated by the range featured in prominent quantum science platforms such as Rydberg atoms ($\alpha=3,6$), trapped ions ($\alpha\in[0,3]$), molecules ($\alpha=3$), and light-matter systems ($\alpha=0$). The limiting case of $\alpha=0$ corresponds to all-to-all interactions, for which the Ising and XX models reduce (up to a constant of motion) to the OAT Hamiltonian $\propto\hat{S}_{z}^{2}$, while f-TAT resolves to the TAT Hamiltonian $\propto(\hat{S}_{x}\hat{S}_{y}+\hat{S}_{y}\hat{S}_{x})$.
II.2 Decoherence
Motivated by the implementation of the above VQC in real-world hardware, we incorporate into our description decoherence due to, for example, stray background fields and spurious couplings to the environment. In our treatment, however, we will make the simplifying assumption that decoherence occurs only during the application of the entangling gate, with the dynamics of the initial and final rotations taken to be sufficiently rapid that they may be approximated as unitary processes.
The inclusion of decoherence and dissipation means that the unitary representation of the entangling gate given by Eq. (1) is inappropriate. Instead, we model this stage of the VQC using a Lindblad master equation,
$\frac{d\hat{\rho}}{dt}=-i[\hat{H}_{\mathrm{I}},\hat{\rho}]+\sum_{\nu,j}\gamma_{\nu}\left(\hat{L}^{\nu}_{j}\hat{\rho}\hat{L}^{\nu\dagger}_{j}-\frac{1}{2}\{\hat{L}^{\nu\dagger}_{j}\hat{L}^{\nu}_{j},\hat{\rho}\}\right),$ (5)
where $\hat{H}_{\mathrm{I}}$ is the chosen entangling Hamiltonian [Eqs. (2)–(4)] and the jump operators $\hat{L}^{\nu}_{k}$ describe decoherence acting on the $k$th qubit at a rate $\gamma_{\nu}$. In this work we focus on two cases: (i) dephasing with $\hat{L}^{\nu}_{k}=\hat{\sigma}_{k}^{\nu}/2$ for $\nu\in\set{x,\>y,\>z}$ and (ii) spontaneous decay and dephasing along $\hat{z}$ given by $\hat{L}^{-}_{k}=\hat{\sigma}^{-}_{k}$ and $\hat{L}^{z}_{k}=\hat{\sigma}_{k}^{z}/2$ at rates $\gamma_{-}$ and $\gamma_{z}$, respectively.
@@ II.3 Optimization of the quantum Fisher information
We seek to identify optimal state preparation protocols, within the constraints of the VQC defined above, for metrology of global spin rotations. The quantity capturing the suitability of a generic quantum state $\hat{\rho}$ for this task is the quantum Fisher information Pang and Brun (2014):
@@
Here |$\ket{k}$ and $\lambda_{k}$ are, respectively, the eigenvectors and corresponding eigenvalues of the density matrix prepared by the VQC; that is, $\hat{\rho}=\sum_{k}\lambda_{k}\ket{k}\!\bra{k}$ |, and $\hat{G}$ is the generator of a unitary transformation $\hat{U}_{\phi}=e^{-i\phi\hat{G}}$ encoding a classical parameter $\phi$. The QFI characterizes the minimum uncertainty with which the classical parameter can be estimated through the quantum Cramer–Rao bound, $(\Delta\phi)^{2}\geq 1/F_{Q}[\hat{\rho}(\mathbf{x});\hat{G}]$. Without loss of generality, throughout this work we set $\hat{G}=\hat{S}_{z}=\sum_{j}\hat{\sigma}^{z}_{j}/2$.
@@
In the following section we instead investigate optimal preparation strategies, using the VQC discussed previously, that take into account the effects of decoherence. For a fixed decoherence strength $\gamma$ we obtain the state $\hat{\rho}(\mathbf{x})$ generated by a given VQC with parameters $\mathbf{x}$. We numerically integrate the master equation (5) for a given choice of entangling Hamiltonian $\hat{H}_{\mathrm{I}}$ and associated global rotations. The optimal preparation strategies defined by parameters $\mathbf{x}_{\mathrm{opt}}$ are obtained by numerically optimizing over $\mathbf{x}$ to maximize the QFI of Eq. (6). Throughout this work we use derivative-free algorithms such as the Nelder–Mead method in SciPy Virtanen et al. (2020) to obtain $\mathbf{x}_{\mathrm{opt}}$. The QFI-Opt package used to simulate the VQC is an open-source Python package Colussi et al. (2024).
III Optimal state preparation with isotropic dephasing